OpenStudy (anonymous):

Solve 4x2 + 2x - 4 = 0 by completing the square. A. (-1 +√17)/4, (-1-√17)/4 B. (1 +√17)/4, (1-√17)/4 C. (5+√37)/2, (5 -√37)/2 D. (-1+√3)/2, (-1-√3)/2

OpenStudy (candyme):

Move the 4 over to the other side. 4x+2x+_=4+_ Half the number with the x and square that 2/2=1 squared= 1 Put that number in the blanks 4x2+2x+1=4+1 4x2+2x+1=5 Now factor

OpenStudy (candyme):

Only this one doesn't factor, so you use the quadratic formula

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

\[4x ^{2}+2x-4=0\] divide by 4, \[or x ^{2}+\frac{ 2 }{4 }x-1=0\] \[x ^{2}+\frac{ 1 }{ 2 }x=1\] adding both sides square of( 1/2)/2,i.e.,1/16 \[x ^{2}+\frac{ 1 }{ 2 }x+\frac{ 1 }{ 16 }=1+\frac{ 1 }{16 }=\frac{ 17 }{1 6 }\] \[\left( x+\frac{ 1 }{ 4 } \right)^{2}=\frac{ 17 }{ 16 }\] \[x+\frac{ 1 }{ 4 }=\pm \frac{ \sqrt{17} }{ 4 }\] \[x+\frac{ 1 }{ 4 }=\frac{ \sqrt{17} }{ 4 },x=\frac{ -1+\sqrt{17} }{ 4 }\] \[x+\frac{ 1 }{ 4 }=\frac{ -\sqrt{17} }{ 4 },x=-\frac{ 1+\sqrt{17} }{ 4 }\]